# On orthogonal decompositions of hermitian Higgs bundles

### v1

@@ 3. Orthogonal decompositions of hermitian Higgs bundles
The aim of this section is to establish extensions of propositions 1 and 4 to the Higgs bundles setting. In particular, we show that it is possible to prove such extended results on Higgs bundles using definitions and results on (holomorphic) Higgs bundles. Hence, strictly speaking, the above extensions only depend on the standard notions given in the previous sections.
According to the terminology introduced in sections 1 and 2 we are now in position to establish the following result.
@@
Let $(\mathfrak{E},h)=(E,\Phi,h)$ be an hermitian Higgs bundle and $D_{h}$, $E^{\prime}$ and $E^{\prime\prime}$ as in Proposition 1 and suppose that $E^{\prime}$ is $\Phi$-invariant. Then both ${\mathfrak{E}}^{\prime}=(E^{\prime},\Phi\lvert_{E^{\prime}})$ and ${\mathfrak{E}}^{\prime\prime}=(E^{\prime\prime},\Phi\lvert_{E^{\prime\prime}})$ are $D_{h}$-invariant Higgs subbundles of ${\mathfrak{E}}=(E,\Phi)$ and they give a Higgs orthogonal decomposition:
@@
From Proposition 1 we know that $E^{\prime},E^{\prime\prime}\subset E$ are $D_{h}$-invariant holomorphic subbundles and we have the holomorphic orthogonal decomposition (1.1), i.e., $E=E^{\prime}\oplus E^{\prime\prime}$ in the holomorphic category.
Since $E^{\prime}$ is holomorphic and also $\Phi$-invariant, the pair ${\mathfrak{E}}^{\prime}=(E^{\prime},\Phi\lvert_{E^{\prime}})$ is a $D_{h}$-invariant Higgs subbundle of ${\mathfrak{E}}=(E,\Phi)$. In particular, by defining ${\mathfrak{S}}={\mathfrak{E}}^{\prime}$ it follows that ${\mathfrak{S}}\subset{\mathfrak{E}}$ is a Higgs subbundle and hence $S^{\bot}=E^{\prime\prime}$ is also $\Phi$-invariant. Since $E^{\prime\prime}$ is holomorphic, it follows that ${\mathfrak{E}}^{\prime\prime}=(E^{\prime\prime},\Phi\lvert_{E^{\prime\prime}})$ is a $D_{h}$-invariant Higgs subbundle of ${\mathfrak{E}}=(E,\Phi)$ and we obtain (3.1). ∎
The above result can be seen as an extension to Higgs bundles of Proposition 1. More precisely, Proposition 7 guarantees that the same classical hypothesis, together with an additional $\Phi$-invariance condition, implies that ${\mathfrak{E}}^{\prime},{\mathfrak{E}}^{\prime\prime}\subset{\mathfrak{E}}$ are Higgs subbundles and that (3.1) is a decomposition of the original Higgs bundle in terms of these Higgs subbundles.
@@
Let $(\mathfrak{E},h)=(E,\Phi,h)$ be an hermitian Higgs bundle, $D_{h}$ the Chern connection of $(E,h)$ and $S\subset E$ a holomorphic subbundle. If the second fundamental form $A_{h}$ of $S$ on $(E,h)$ vanishes identically and $S$ is $\Phi$-invariant, then ${\mathfrak{S}}=(S,\Phi\lvert_{S})$ and ${\mathfrak{S}}^{\bot}=(S^{\bot},\Phi\lvert_{S^{\bot}})$ are Higgs subbundles of ${\mathfrak{E}}=(E,\Phi)$ and
@@
From Proposition 4 we know that $S^{\bot}\subset E$ is a holomorphic subbundle and we have the holomorphic orthogonal decomposition (1.8), i.e., $E=S\oplus S^{\bot}$ in the holomorphic category.
Since $S\subset E$ is holomorphic and $\Phi$-invariant, the pair ${\mathfrak{S}}=(S,\Phi\lvert_{S})$ is a Higgs subbundle of ${\mathfrak{E}}=(E,\Phi)$. In particular, $S^{\bot}$ is $\Phi$-invariant and since it is also holomorphic, the pair ${\mathfrak{S}}^{\bot}=(S^{\bot},\Phi\lvert_{S^{\bot}})$ is a Higgs subbundle of ${\mathfrak{E}}=(E,\Phi)$ and we obtain (3.2). ∎
The above proposition can be seen as a Higgs extension of Proposition 4. In particular, Proposition 8 guarantees that the same classical hypothesis, together with an additional $\Phi$-invariance condition, implies that ${\mathfrak{S}},{\mathfrak{S}}^{\bot}\subset{\mathfrak{E}}$ are Higgs subbundles and that (3.2) is a decomposition of the original Higgs bundle in terms of these Higgs subbundles. It is important to mention that this proposition can also be proved using Proposition 7 and Lemma 2. In fact, by redefining $E^{\prime}=S$ and $E^{\prime\prime}=S^{\bot}$ and since $E^{\prime}$ is $\Phi$-invariant, then ${\mathfrak{E}}^{\prime}={\mathfrak{S}}\subset{\mathfrak{E}}$ is a Higgs subbundle. If $A_{h}$ vanishes identically, (2.5) implies that $E^{\prime}$ is ${\cal D}_{h}$-invariant. Therefore, from Lemma 2 we know that $E^{\prime}$ is also $D_{h}$-invariant and the result follows from Proposition 7.
As in the classical case, the converse of Proposition 8 is also true. In fact, if ${\mathfrak{E}}={\mathfrak{S}}\oplus{\mathfrak{S}}^{\bot}$ is a Higgs decomposition, the pairs $({\mathfrak{S}},h_{S})$ and $({\mathfrak{S}}^{\bot},h_{S^{\bot}})$ are both hermitian Higgs bundles. If ${\cal D}_{h_{S}}$ and ${\cal D}_{h_{S^{\bot}}}$ are the Hitchin-Simpson connections of these pairs, then ${\cal D}_{h}$ is the direct sum of ${\cal D}_{h_{S}}$ and ${\cal D}_{h_{S^{\bot}}}$ and the result follows from (2.5) and Proposition 5, or equivalently from (2.7) and Proposition 6.
We are now in position to mention some applications of the above results in the Higgs bundles setting. Recall that in Section 2 we have mentioned that the second fundamental form of a Higgs subbundle of a hermitian Higgs bundle is the same as the second fundamental form of the corresponding holomorphic subbundle. From this fact, the previous paragraph and Proposition 8, we immediately get the following result.
@@
Let $(\mathfrak{E},h)=(E,\Phi,h)$ be an hermitian Higgs bundle, ${\cal D}_{h}$ the Hitchin-Simpson connection of $({\mathfrak{E}},h)$ and ${\mathfrak{S}}\subset{\mathfrak{E}}$ a Higgs subbundle. Then, the second fundamental form $A_{h}$ of ${\mathfrak{S}}$ in $({\mathfrak{E}},h)$ given by (2.5) vanishes identically if and only if ${\mathfrak{E}}={\mathfrak{S}}\oplus{\mathfrak{S}}^{\bot}$ is a Higgs orthogonal decomposition.
We would like to emphasize that this result can be completely established in terms of objects in the Higgs category, therefore it is natural to wonder if there exists another proposition, closely related to Proposition 7, and such that can be established making reference only to objects in the Higgs category. As we will see in a moment, it is indeed the case and Lemma 2 is a key result in the proof of such a proposition. Before doing that, let us introduce the following definition. Let ${\mathfrak{E}}^{\prime}\subset{\mathfrak{E}}$ be a Higgs subbundle and $E^{\prime\prime}\subset E$ the orthogonal complement of $E^{\prime}$ with respect to $h$. Let us consider the pair ${\mathfrak{E}}^{\prime\prime}=(E^{\prime\prime},\Phi\lvert_{E^{\prime\prime}})$, which we call the $C^{\infty}$ Higgs orthogonal complement of the Higgs subbundle ${\mathfrak{E}}^{\prime}$ with respect to $h$. The bundle $E^{\prime\prime}\subset E$ is just a $C^{\infty}$ complex subbundle (not necessarily holomorphic) and hence, strictly speaking, ${\mathfrak{E}}^{\prime\prime}$ is not necessarily a Higgs subbundle of ${\mathfrak{E}}$. However, that will be the case if certain invariant condition holds. To be precise, we have the following result.
@@
Let $(\mathfrak{E},h)=(E,\Phi,h)$ be an hermitian Higgs bundle and ${\cal D}_{h}$ the Hitchin-Simpson connection of $({\mathfrak{E}},h)$. Let ${\mathfrak{E}}^{\prime}\subset{\mathfrak{E}}$ be a ${\cal D}_{h}$-invariant Higgs subbundle and ${\mathfrak{E}}^{\prime\prime}$ the $C^{\infty}$ Higgs orthogonal complement of ${\mathfrak{E}}^{\prime}$ with respect to $h$. Then ${\mathfrak{E}}^{\prime\prime}\subset{\mathfrak{E}}$ is also a ${\cal D}_{h}$-invariant Higgs subbundle and ${\mathfrak{E}}={\mathfrak{E}}^{\prime}\oplus{\mathfrak{E}}^{\prime\prime}$ is a Higgs orthogonal decomposition.
@@
By applying Lemma 2 to the Higgs subbundle ${\mathfrak{E}}^{\prime}\subset{\mathfrak{E}}$ we know that $E^{\prime}$ is $D_{h}$-invariant. Since it is also $\Phi$-invariant, Proposition 7 implies the Higgs orthogonal decomposition ${\mathfrak{E}}={\mathfrak{E}}^{\prime}\oplus{\mathfrak{E}}^{\prime\prime}$ with ${\mathfrak{E}}^{\prime\prime}\subset{\mathfrak{E}}$ also a $D_{h}$-invariant Higgs subbundle. The ${\cal D}_{h}$-invariance of this last subbundle follows again from Lemma 2. ∎

### v2

@@ 4. Orthogonal decompositions of hermitian Higgs bundles
The aim of this section is to establish extensions of propositions 1 and 4 to the Higgs bundles setting. In particular, we show that these extended results on Higgs bundles can be obtained from the usual definitions and results on (holomorphic) Higgs bundles if additional invariance conditions hold. Hence, strictly speaking, the above extensions rely only on the standard notions given in the previous sections.
Let $\mathcal{S}\subset\mathfrak{E}$ be a Higgs subbundle. In order to extend propositions 1 and 4 lets asume that $\mathcal{S}$ is also ${\bar{\Phi}}_{h}$-invariant. Notice that this condition is equivalent to $S^{\perp}$ being $\Phi$-invariant. From the above we immediately get the following result.
Lemma 2.
If $\mathfrak{S}\subset\mathfrak{E}$ is a Higgs subbundle that is ${\bar{\Phi}}_{h}$-invariant, then the holomorphic subbundle $S\subset E$ is $D_{h}$-invariant if and only if it is $\mathcal{D}_{h}$-invariant.
According to the terminology introduced in sections 2 and 3, we are now in a position to establish the following result.
@@
Let $(\mathfrak{E},h)=(E,\Phi,h)$ be a hermitian Higgs bundle and $D_{h}$, $E^{\prime}$ and $E^{\prime\prime}$ as in Proposition 1 and suppose that $E^{\prime}$ is $\Phi$-invariant and also ${\bar{\Phi}}_{h}$-invariant. Then both ${\mathfrak{E}}^{\prime}=(E^{\prime},\Phi\lvert_{E^{\prime}})$ and ${\mathfrak{E}}^{\prime\prime}=(E^{\prime\prime},\Phi\lvert_{E^{\prime\prime}})$ are $D_{h}$-invariant Higgs subbundles of ${\mathfrak{E}}=(E,\Phi)$ and they give a Higgs orthogonal decomposition:
@@
From Proposition 1 we know that $E^{\prime},E^{\prime\prime}\subset E$ are $D_{h}$-invariant holomorphic subbundles and we have the holomorphic orthogonal decomposition (2.2), i.e., $E=E^{\prime}\oplus E^{\prime\prime}$ in the holomorphic category. Since $E^{\prime}$ is holomorphic and also $\Phi$-invariant, the pair ${\mathfrak{E}}^{\prime}=(E^{\prime},\Phi\lvert_{E^{\prime}})$ is a $D_{h}$-invariant Higgs subbundle of ${\mathfrak{E}}=(E,\Phi)$. In particular, by defining ${\mathfrak{S}}={\mathfrak{E}}^{\prime}$ it follows that ${\mathfrak{S}}\subset{\mathfrak{E}}$ is a Higgs subbundle and hence $S^{\bot}=E^{\prime\prime}$ is also $\Phi$-invariant. Since $E^{\prime\prime}$ is holomorphic, it follows that ${\mathfrak{E}}^{\prime\prime}=(E^{\prime\prime},\Phi\lvert_{E^{\prime\prime}})$ is a $D_{h}$-invariant Higgs subbundle of ${\mathfrak{E}}=(E,\Phi)$ and we obtain (4.1). ∎
The above result can be seen as an extension to Higgs bundles of Proposition 1. More precisely, Proposition 7 guarantees that the same classical hypothesis, together with additional $\Phi$ and ${\bar{\Phi}}_{h}$-invariance conditions, implies that ${\mathfrak{E}}^{\prime},{\mathfrak{E}}^{\prime\prime}\subset{\mathfrak{E}}$ are Higgs subbundles and that (4.1) is a decomposition of the original Higgs bundle in terms of these Higgs subbundles. In particular, let $(\mathfrak{E},h)=(E,\Phi,h)$ be a hermitian Higgs bundle and let us consider again the decomposition (2.4) for the holomorphic bundle $E$. If $E^{0},...,E^{k}$ are all $\Phi$ and ${\bar{\Phi}}_{h}$-invariant and $\Phi_{0},\Phi_{1},...,\Phi_{k}$ denote the restrictions of $\Phi$ to these bundles, respectively, Proposition 7 immediately implies that ${\mathfrak{E}}^{0}=(E^{0},\Phi_{0}),{\mathfrak{E}}^{1}=(E^{1},\Phi_{1}),...,{\mathfrak{E}}^{k}=(E^{k},\Phi_{k})$ are all Higgs subbundles of the original Higgs bundle $\mathfrak{E}$}. To be precise, in such a case
(4.2) ${\mathfrak{E}}={\mathfrak{E}}^{0}\oplus{\mathfrak{E}}^{1}\oplus\cdots\oplus{\mathfrak{E}}^{k}\,.$
is a Higgs orthogonal decomposition of $\mathfrak{E}$}.
@@
Let $(\mathfrak{E},h)=(E,\Phi,h)$ be a hermitian Higgs bundle, $D_{h}$ the Chern connection of $(E,h)$ and $S\subset E$ a holomorphic subbundle. If the second fundamental form $A_{h}$ of $S$ on $(E,h)$ vanishes identically and $S$ is $\Phi$ and ${\bar{\Phi}}_{h}$-invariant, then ${\mathfrak{S}}=(S,\Phi\lvert_{S})$ and ${\mathfrak{S}}^{\bot}=(S^{\bot},\Phi\lvert_{S^{\bot}})$ are Higgs subbundles of ${\mathfrak{E}}=(E,\Phi)$ and
@@
From Proposition 4 we know that $S^{\bot}\subset E$ is a holomorphic subbundle and we have the holomorphic orthogonal decomposition (2.11), i.e., $E=S\oplus S^{\bot}$ in the holomorphic category. Since $S\subset E$ is holomorphic and $\Phi$-invariant, the pair ${\mathfrak{S}}=(S,\Phi\lvert_{S})$ is a Higgs subbundle of ${\mathfrak{E}}=(E,\Phi)$. In particular, $S^{\bot}$ is $\Phi$-invariant and since it is also holomorphic, the pair ${\mathfrak{S}}^{\bot}=(S^{\bot},\Phi\lvert_{S^{\bot}})$ is a Higgs subbundle of ${\mathfrak{E}}=(E,\Phi)$ and we obtain (4.3). ∎
The above proposition can be seen as a Higgs extension of Proposition 4. In particular, Proposition 8 guarantees that the same classical hypothesis, together with additional $\Phi$ and ${\bar{\Phi}}_{h}$-invariance conditions, implies that ${\mathfrak{S}},{\mathfrak{S}}^{\bot}\subset{\mathfrak{E}}$ are Higgs subbundles and that (4.3) is a decomposition of the original Higgs bundle in terms of these Higgs subbundles. As in the classical case, the converse of Proposition 8 is also true. In fact, if ${\mathfrak{E}}={\mathfrak{S}}\oplus{\mathfrak{S}}^{\bot}$ is a Higgs decomposition, the pairs $({\mathfrak{S}},h_{S})$ and $({\mathfrak{S}}^{\bot},h_{S^{\bot}})$ are both hermitian Higgs bundles. If ${\mathcal{D}}_{h_{S}}$ and ${\mathcal{D}}_{h_{S^{\bot}}}$ are the Hitchin-Simpson connections of these pairs, then ${\mathcal{D}}_{h}$ is the direct sum of ${\mathcal{D}}_{h_{S}}$ and ${\mathcal{D}}_{h_{S^{\bot}}}$ and the result follows from (3.6), (3.7) and Proposition 5.
We are now in a position to mention some applications of the above results in the Higgs bundles setting. Recall that in Section 3 we have mentioned that the form ${\mathcal{A}}_{h}$ is the same as the second fundamental form $A_{h}$ of the corresponding holomorphic subbundle, provided the Higgs bundle satisfies an additional ${\bar{\Phi}}_{h}$-invariance condition. From this fact, the previous paragraph and Proposition 8, we immediately get the following result.
@@
Let $(\mathfrak{E},h)=(E,\Phi,h)$ be a hermitian Higgs bundle and ${\mathcal{D}}_{h}$ its Hitchin-Simpson connection. Let ${\mathfrak{S}}\subset{\mathfrak{E}}$ be a Higgs subbundle that is ${\bar{\Phi}}_{h}$-invariant. Then, the second fundamental form $\mathcal{A}_{h}$ of $\mathfrak{S}$} in $({\mathfrak{E}},h)$ given by (3.7) vanishes identically if and only if ${\mathfrak{E}}={\mathfrak{S}}\oplus{\mathfrak{S}}^{\bot}$ is a Higgs orthogonal decomposition.
We would like to emphasize that this result can be completely established in terms of objects in the Higgs category. Therefore, it is natural to wonder if there exists another proposition, closely related to Proposition 7, and which can be established using only objects in the Higgs category. As we will see in a moment, it is indeed the case and Lemma 2 is a key result in the proof of such a proposition. Before doing that, let us introduce the following definition. Let ${\mathfrak{E}}^{\prime}\subset{\mathfrak{E}}$ be a Higgs subbundle and $E^{\prime\prime}\subset E$ the orthogonal complement of $E^{\prime}$ with respect to $h$. Let us consider the pair ${\mathfrak{E}}^{\prime\prime}=(E^{\prime\prime},\Phi\lvert_{E^{\prime\prime}})$, which we call the $C^{\infty}$ Higgs orthogonal complement of the Higgs subbundle $\mathfrak{E}$$^{\prime}$ with respect to $h$. The bundle $E^{\prime\prime}\subset E$ is just a $C^{\infty}$ complex subbundle (not necessarily holomorphic) and hence, strictly speaking, $\mathfrak{E}$$^{\prime\prime}$ is not necessarily a Higgs subbundle of $\mathfrak{E}$}. However, that will be the case if a certain invariant condition holds. To be precise, we have the following result.
@@
Let $(\mathfrak{E},h)=(E,\Phi,h)$ be a hermitian Higgs bundle and ${\mathcal{D}}_{h}$ its Hitchin-Simpson connection. Let ${\mathfrak{E}}^{\prime}\subset{\mathfrak{E}}$ be a ${\mathcal{D}}_{h}$ and ${\bar{\Phi}}_{h}$-invariant Higgs subbundle and $\mathfrak{E}$$^{\prime\prime}$ the $C^{\infty}$ Higgs orthogonal complement of $\mathfrak{E}$$^{\prime}$ with respect to $h$. Then ${\mathfrak{E}}^{\prime\prime}\subset{\mathfrak{E}}$ is also a ${\mathcal{D}}_{h}$-invariant Higgs subbundle and ${\mathfrak{E}}={\mathfrak{E}}^{\prime}\oplus{\mathfrak{E}}^{\prime\prime}$ is a Higgs orthogonal decomposition.
@@
By applying Lemma 2 to the Higgs subbundle ${\mathfrak{E}}^{\prime}\subset{\mathfrak{E}}$ we know that $E^{\prime}$ is $D_{h}$-invariant. Since it is also $\Phi$-invariant, Proposition 7 implies the Higgs orthogonal decomposition ${\mathfrak{E}}={\mathfrak{E}}^{\prime}\oplus{\mathfrak{E}}^{\prime\prime}$ with ${\mathfrak{E}}^{\prime\prime}\subset{\mathfrak{E}}$ also a $D_{h}$-invariant Higgs subbundle. The ${\mathcal{D}}_{h}$-invariance of this last subbundle follows again from Lemma 2. ∎
5. The Kobayashi functional for Higgs bundles
Let ${\mathfrak{E}}=(E,\Phi)$ be a Higgs bundle and $h$ a hermitian metric in $\mathfrak{E}$}. If ${\mathcal{R}}_{h}$ is the Hitchin-Simpson curvature defined in (3.2), we can define the Hitchin-Simpson mean curvature ${\mathcal{K}}_{h}$ as the element in $A^{0}({\rm End}E)$ satisfying
(5.1) $in{\mathcal{R}}_{h}\wedge\omega^{n-1}={\mathcal{K}}_{h}\,\omega^{n}\,.$
The above definition is the natural extension to Higgs bundles of the classical definition (2.1) for holomorphic vector bundles. A hermitian metric $h$ in $\mathfrak{E}$} is called a Hermitian-Yang-Mills metric if ${\mathcal{K}}_{h}=cI$, where $c$ is the same constant introduced in Section 2. If $h$ is Hermitian-Yang-Mills, the pair $({\mathfrak{E}},h)$ is called a Hermitian-Yang-Mills Higgs bundle.
Let us denote by ${\rm Herm}^{+}{\mathfrak{E}}$ the space of hermitian metrics [^9] It is clearly the same space ${\rm Herm}^{+}E$ defined in Section 2, the change of notation is just to emphasize that we are now interested in the Higgs bundles setting. in the Higgs bundle $\mathfrak{E}$}. If $h\in{\rm Herm}^{+}{\mathfrak{E}}$ is considered as a variable and ${\mathcal{K}}_{h}$ is the Hitchin-Simpson curvature defined in (5.1), then we can consider the functional ${\mathcal{J}}:{\rm Herm}^{+}{\mathfrak{E}}\longrightarrow{\mathbb{R}}$ defined by
(5.2) ${\mathcal{J}}(h)=\frac{1}{2}\int_{M}\lvert{\mathcal{K}}_{h}\lvert^{2}\omega^{n}=\frac{n!}{2}\lVert{\mathcal{K}}_{h}\lVert^{2}\,$
which is, up to a multiplicative constant, the natural energy functional for the Hitchin-Simpson mean curvature. The functional (5.2) is called the Kobayashi functional for the Higgs bundle $\mathfrak{E}$} and is the obvious extension to Higgs bundles of the classical functional (2.12) for the corresponding holomorphic vector bundle $E$. It can be shown (see [6] for details) that a hermitian metric $h$ in $\mathfrak{E}$} is Hermitian-Yang-Mills if and only if it is a minimum of (5.2) and more in general, one has the following result.
Theorem 3.
([6], Thm. 4). Let $\mathfrak{E}$} be a Higgs bundle. A hermitian metric $h$ is a critical point of the functional (5.2) if and only if the Hitchin-Simpson mean curvature is parallel with respect to the Hitchin-Simpson connection defined by $h$, i.e., if and only if
(5.3) ${\mathcal{D}}_{h}{\mathcal{K}}_{h}=0\,.$
Notice that Theorem 3 is the generalization to Higgs bundles of Theorem 1 for holomorphic vector bundles. On the other hand, the Hitchin-Simpson connection ${\mathcal{D}}_{h}$ is neither compatible with the holomorphic nor the hermitian structure of $E$, a fact also mentioned in [2, 16]. In fact, the (0,1) part of ${\mathcal{D}}_{h}$ is $d^{\prime\prime}+\bar{\Phi}_{h}$ which coincides with $d^{\prime\prime}$ if and only if $\Phi$ is identically zero. Moreover, just from the definition of the Hitchin-Simpson connection one has
(5.4) ${\mathcal{D}}_{h}h=(D_{h}+\Phi+\bar{\Phi}_{h})h=(\Phi+\bar{\Phi}_{h})h\,.$
Here we have used the fact that $h$ is parallel with respect to the Chern connection $D_{h}$. We should notice that the right hand side in the above expression does not vanish [^10] Let $s=(s_{i})$ be a local frame for $E$ and $s^{*}=(s^{j})$ its dual frame. Then $h=\sum h_{j{\bar{k}}}s^{j}\otimes{\bar{s}}^{k}$ and for any connection $D$ in $E$ one has $Dh=\sum(dh_{i{\bar{\jmath}}}-h_{k{\bar{\jmath}}}\,\theta^{k}_{i}-h_{i{\bar{k}}}\,{\bar{\theta}}^{k}_{j})s^{i}\otimes{\bar{s}}^{j}$ where $\theta=(\theta^{k}_{i})$ is the connection form of $D$ (see [13], p. 18 for details). Since $D_{h}h=0$ one has (5.5) ${\mathcal{D}}_{h}h=(\Phi+\bar{\Phi}_{h})h=-\sum(h_{k{\bar{\jmath}}}\,\phi^{k}_{i}+h_{i{\bar{k}}}\,{\bar{\phi}}^{k}_{j}+h_{k{\bar{\jmath}}}(\phi^{*}_{h})^{k}_{i}+h_{i{\bar{k}}}{\overline{(\phi^{*}_{h})}}^{k}_{j})s^{i}\otimes{\bar{s}}^{j},$ where $\phi^{k}_{i}$ and $({\phi^{*}_{h}})_{i}^{k}$ are the components of $\Phi$ and $\bar{\Phi}_{h}$. Now, ${\mathcal{D}}_{h}h=0$ if and only if its $(1,0)$ and $(0,1)$ parts are both equal to zero. In particular, if $s$ is unitary the $(1,0)$ part is $\sum h_{k{\bar{\jmath}}}\,\phi^{k}_{i}-\sum h_{i{\bar{k}}}{\overline{(\phi^{*}_{h})}}^{k}_{j}=-\phi^{j}_{i}-{\overline{(\phi^{*}_{h})}}^{i}_{j}=-2\phi^{j}_{i}$ which does not vanish unless $\Phi=0$. and hence (5.4) establishes a crucial fact. Namely, for any proper Higgs bundle (i.e., one in which $\Phi$ is not identically zero), $h$ is not parallel with respect to the Hitchin-Simpson connection ${\mathcal{D}}_{h}$. As a consequence of this, it is not possible to obtain a straightforward adaptation to Higgs bundles of the classical arguments of Section 2 where, as we have shown before, the parallelism condition of $h$ plays a key role. In fact, the lack of a parallelism condition of the metric $h$ with respect to ${\mathcal{D}}_{h}$ shows a significant difference between the Higgs bundles setting and the classical case of holomorphic vector bundles. It is just because of the non-vanishing of (5.4) that it is not possible to generalize to Higgs bundles the classical Theorem 2, which gives a classification of the critical points of the Kobayashi functional. This makes it explicit, once again, that the functional $\mathcal{J}$} defined in (5.2) does not always satisfy the same properties of the classical counterpart $J$, introduced in (2.12). Indeed, similar contrasts between the Higgs bundles setting and the classical one have already been mentioned in [6], where it was shown that the difference between $\mathcal{J}$} and the Yang-Mills-Higgs functional is given by a couple of terms, one of these a topological constant involving the first and second Chern classes of $\mathfrak{E}$}, and the other one a term depending on the hermitian metric. A situation in clear contrast with the classical case, where the Kobayashi and the Yang-Mills functional differ only by the topological constant.
6. Discussion and conclusions
In general terms, the central object of this article was the notion of Higgs bundle. As previously mentioned, Higgs bundles arose as geometric objects associated with the two-dimensional reduction of the self-dual Yang–Mills equations in four dimensions. Consequently, all properties of Higgs bundles are potentially relevant not only in complex geometry, but also in Yang-Mills theory, and more in general in quantum field theory and string theory. In order to make this article accessible to a broader audience, including both geometers and mathematical physicists, we have included a Section 2 in the form of preliminaries, where we summarized the main classical material from complex geometry that will be important in the rest of the article. In this section, we chose to follow the notation of the classical text of Kobayashi [13], as it provided an elegant and aesthetically pleasing framework that is also well suited to Higgs bundles. The aim of this section was to bring together several properties that appeared in different chapters of [13], but which are often scattered throughout other classical texts on complex geometry, such as those of Demailly [8] and Griffiths and Harris [9]. Additionally, we firmly believe that this indeed makes the article more self-contained. Also, in this section, we presented the key ideas without relying on local computations, and, as we saw, this strategy proved useful throughout the rest of the paper. It is worth noting that a similar treatment—avoiding local computations—was also employed in [15], although in the context of the of Spin bundles.
As we mentioned in the first two sections, the main purpose of this article has been to study certain geometric properties of Higgs bundles. In particular, we focused on results that can be seen as natural extensions of classical results in complex geometry to the context of Higgs bundles. In Section 3 we have shown that if $({\mathfrak{E}},h)=(E,\Phi,h)$ is a hermitian Higgs bundle and ${\mathfrak{S}}\subset{\mathfrak{E}}$ is a Higgs subbundle, then $S^{\perp}$ is ${\bar{\Phi}}_{h}$-invariant, but not necessarily $\Phi$-invariant. This fact played a crucial role and was the origin of the forms ${\mathcal{A}}_{h}$ and ${\mathcal{B}}_{h}$ that appeared in the Gauss-Codazzi equations for hermitian Higgs bundles. Here, it is important to note that, even though the form of the Gauss–Codazzi equations (3.13) was formally similar to the classical ones, the new forms ${\mathcal{A}}_{h}$ and ${\mathcal{B}}_{h}$ were not the same as $A_{h}$ and $B_{h}$. A fact that followed from the definitions (3.6) and (3.10). In Section 4 we have shown that using additional invariance conditions we obtain some propositions in the Higgs bundle settings that can be considered as natural extensions of classical results on holomorphic hermitian vector bundles. In particular, at the end of this section we obtain a couple of corollaries that can be established only in terms of objects in the Higgs category.
Finally, as a straightforward consequence of Theorem 3 of Section 5, if $h$ is a critical point of the Kobayashi functional $\mathcal{J}$}, then the curvature ${\mathcal{K}}_{h}$ satisfies a parallelism condition with respect to the Hitchin-Simpson connection ${\mathcal{D}}_{h}$. However, the metric $h$ is not parallel with respect to ${\mathcal{D}}_{h}$, a fact that is summarized in (5.4). As a consequence of this, we have shown that, as far as the Kobayashi functional is concerned, a classical decomposition theorem cannot be naturally extended to Higgs bundles. We do not know yet if there exists a kind of decomposition result concerning the critical points of the functional (5.2). We hope to address this question in a subsequent article.
Acknowledgements
The first author wants to thank H. García-Compeán for bringing Wijnholt’s paper [22] for their attention. The second author wants to thank the support of Secretaría de Ciencia, Humanidades, Tecnología e Innovación (Secihti) through the doctoral scholarship No. 830628.